OpenStudy (anonymous):

u(0,t) = u (1,t) = 0 u(x,0) = 0,01sen3πx du/dt (x,0) = 0 help!

OpenStudy (anonymous):

are you trying to write $$u(0,t)=u(1,t)=0\\u(x,0)=0.01\sin(3\pi x)\\\frac{du}{dt}\Biggr|_{t=0}=0$$

OpenStudy (anonymous):

what equation are these boundary conditions for?

OpenStudy (anonymous):

means that g (x) = 0

OpenStudy (anonymous):

eh... :-( I don't follow, sorry. you haven't stated anything

OpenStudy (anonymous):

what you gave are just boundary conditions for some PDE presumably... maybe heat equation?

OpenStudy (anonymous):

wave equation

OpenStudy (anonymous):

\[u ( x, t ) = (sen n pix/l)\left[ an \cos (npict/l) \right]\]

OpenStudy (anonymous):

an=2 \[\int\limits_{0}^{1}0.01sen3pix \times sen n \pi x dx \]

OpenStudy (anonymous):

so you're saying \(u(x,t)=\sum a_n\sin(n\pi x)\cos(n\pi ct/l)\)?

OpenStudy (anonymous):

oops \(\sin(n\pi x/l)\)

OpenStudy (anonymous):

yes